OpenStudy (anonymous):

I dont understand sequencing and how to do it. the sequence is 2 6 3 8 6 12. what is the next number. how do I figure this out.

OpenStudy (amistre64):

2+4 = 6 6-3 = 3 3+5 = 8 8-2 = 6 6+6 = 12 just try to find a pattern?

OpenStudy (amistre64):

+4.+5 +6 on every other one -3 -2 -1 next?

OpenStudy (amistre64):

11 is my guess

OpenStudy (anonymous):

thanks, I am not getting the patterns figured out very well.

OpenStudy (amistre64):

trial and error really :)

OpenStudy (anonymous):

11 is right.

OpenStudy (anonymous):

+4 -3.+5 -2 +6 Then -1 next would make it 11.